How do you work out 3(y+2)=24
Answer by queenofit(120)

First you must distribute you 3 throughout the portion of the equation that is inside the ( ). this is done by multiplication. multiply the 3 to the y and then to the 2 to set up the problem like this.
3(y+2)=24
3*y=3y
3*2=6
so now you have the equation
3y+6=24
you must get the numbers and the variables on different sides of the = sign so move the 6 to the other side by subtracting 6 from both sides. This gets rid of the six on the left side of your equal sign. And you will have
3y=24-6
3y=18
but you still have a 3y and you are solving for y. so you must divide each side by the three. Remember, anything you do to one side of the = you must do to the other side of the = . so you have
3y/3=18/3
y=6
Remember to watch your +/- signs when you are working any equations. This is where most people make their mistakes. Remember what you do to one side you must do to the other side.
hope this helps.